OpenStudy (anonymous):

Find the slope of the line that passes through the points (-2, 4) and (-5, -6). -2/7 10/3 is my answer, am I correct? -2/3

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (-2,4) x1 = -2 and y1 = 4 (-5,-6) x2 = -5 and y2 = -6 now we sub slope = (-6-4) / (-5-(-2)) slope = -10/(-5 + 2) slope = -10/-3 = 10/3 you are correct :)

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

no problem...you knew what you were doing :)

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

Find the slope of the line whose equation is 8y = 4 - 3x. -3/8 -8/3 2

OpenStudy (texaschic101):

you need to put it in y = mx + b form where m is your slope 8y = -3x + 4 -- divide by 8 y = -3/8x + 1/2 slope is -3/8

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)